OpenStudy (anonymous):

help plz y=x^2+2x y=3x+20

OpenStudy (faiqraees):

Substitute the second equation into first equation

OpenStudy (anonymous):

the question is solve the system of equations algebraically show all of your steps

OpenStudy (faiqraees):

yeah I know

OpenStudy (faiqraees):

y=x²+2x y = 3x +20 x²+2x = y = 3x+20 x²+2x = 3x+20 Now solve for x

OpenStudy (anonymous):

x=-4